OpenStudy (anonymous):

A quadrilateral contains two equal sides measuring 12 cm each with an included right angle. If the measure of the third side is 8 cm and the angle opposite the right angle is 120 degrees, find the fourth side and the area of the quadrlateral.

OpenStudy (anonymous):

draw it: make a-b-c be right angle with sides 12cm long and 90deg between them make d be point that is somewhere between a and c but 8cm from c. distance abc is right angle triangle with hypotenuse ac=12*sqrt(2). then using cosine law we find missing side. note that there will be two solutions: b^2 = a^2 + c^2 - 2ac*cos(120)

OpenStudy (anonymous):

I can't draw it that's the problem =.=

OpenStudy (anonymous):

The corresponding diagonal of the quadr. is rt288. Now let fourth side be x. Use cosine rule and you get. 288 = 64 + x^2 - 2(8)cos120. Which leads you to x = 4+16rt15. Now use sine rule to find the area of each triangle and add them.

OpenStudy (anonymous):

fig please =.=

Directrix (directrix):

Here is the problem or a similar as worked here a few days ago. http://openstudy.com/study?#/updates/5275f0c5e4b0f6d98a1f7ae9